# Effective Dynamics of Spherically Symmetric Static Spacetime

### v1

@@ V.3 Generating Kawai-Yokokura metrics
at leading orders for large mass, $a_{0}\gg l_{p}$. Note that we write here $r_{s}$ for the surface radius, instead of the Schwarzschild radius. Indeed, it must be underlined that $r_{s}>a_{0}$ and that there is no actual horizon555 Note that the proper length between $r=r_{s}$ and $r=a_{0}$ is of order $\sqrt{n}l_{p}\gg l_{p}$. . Phenomenologically, however, the exponentially large redshift of the KY metric (87) makes the imaging almost black and, seen from the exterior, the object behaves as a black hole in astrophysical cases with non-Planckian masses. The small deviation from classical Schwarzschild black holes is consistent with current observations CY .
@@
Indeed, it was shown in KY4 that the KY metric (87) is a non-perturbative solution in $\hbar$ for $\eta\in[1,2]$ and $\sigma=\mathcal{O}(nl_{p}^{2})$, where $n$ is the number of the degrees of freedom that contribute to the entropy of the system. More precisely, for $n$ massless scalar fields minimally-coupled to gravity, we have KY4 :
@@
The KY metric leads to a couple of interesting physical insights. First, it can be obtained from several paths. The original way KMY is to study the time evolution of a spherical collapsing matter including the backreaction from the particle creation in the time-dependent spacetime according to (89). The final object at equilibrium in a heat bath is described by the KY metric666 There exist similar constructions incorporating various aspects of QFT fluctuations in the structure of black holes. A successful example are black stars that include the backreaction of vacuum polarization at every point inside the compact object (through a renormalized stress-energy tensor), creating an onion structure with Hawking radiation between each shell generating a negative pressure balancing out the gravitational pull Barcelo:2007yk ; Visser:2008rtf ; Barcelo:2009tpa . The dynamical structure of KY black holes is very similar, and further incorporates non-perturbative 4D quantum fluctuations (beyond the vacuum state) through a detailed balance of the 4D anomaly. The Hawking radiation within the compact object is a mere semi-classical expression of those quantum fluctuations. KY1 ; KY2 ; KY3 ; KY4 ; KY5 . Another path is to identify the KY metric as the entropy-maximizing configuration for a given surface area $\mathcal{A}$, saturating the Bousso bound Y2 .
@@
for $l_{p}\ll r\leq r_{s}$, where the constant factor $b_{0}=\frac{\eta}{2}e^{-\frac{r_{s}^{2}}{4\eta\sigma}}$ encodes the mass (and thus the surface radius) of the black hole. Such an effective Hamiltonian is supposed to reflect the full non-perturbative coupling of quantum matter fields on the geometry (such as the curvature generated by the one-loop renormalized QFT corrections).
@@
We can check that solving the corresponding equations of motion for $a$ and $b$, (66) and (67), leads back to the desired behavior for $r\gg l_{p}$:
@@
Solving exactly the equation of motion for $a$ leads to the desired behavior,
@@
[…]
@@
[…]
where $C$ and $r_{0}$ are two integration constants777 The extra integration constant $C$ does not affect the large scale behavior of the metric at $r\gg l_{P}$, but it suggests various versions and completions of the KY metric to the central region $0\leq r\lesssim l_{p}$. . Plugging the explicit expressions for $c_{1}$ and $c_{3}$ yields:
[…]
@@
This shows that our type-$(+1)$ ansatz for effective Hamiltonians, linear in the lapse $b$ and polynomial in the mass function $a$, is rich enough to contain a whole family of Hamiltonians generating the KY metric for $r\gg l_{p}$. More precisely, the expression of the effective energy density $\rho_{\mathrm{eff}}(a,r)$ is not unique and there are various expressions of $\rho_{\mathrm{eff}}(a,r)$ that reproduce the same KY metric. On the one hand, having multiple effective Hamiltonian leading to the KY metric underlines a certain universality of this solution. On the other hand, it shows that certain (combinations of the) coefficients $c_{n}$ in the polynomial expansion of $\rho_{\mathrm{eff}}$ are relevant to semi-classical behavior of the space-time metric (for $r\gg l_{P}$), while others are only relevant to the microscopic regime for the Planck scale physics at $0\leq r\lesssim l_{p}$.
Then, those various admissible Hamiltonians a priori predict different physics and geometries for the central region $0\leq r\lesssim l_{p}$. The completion of the KY space-time to the whole black hole interior will therefore be a strong criteria for selecting appropriate effective Hamiltonians. As shown in the next section, this will allow us to identify a fully regular completion of the KY metric, without singularity for the whole interior region $0\leq r\leq r_{s}$.
@@ VI.1 Conditions on the Energy Density
We first examine the conditions on $a(r)$ and $b(r)$ for the spherically-symmetric static metric (3) to be regular at $r=0$. Checking the behavior of the curvatures as $r\to 0$, as proven in details in appendix A, one can show that the necessary behavior for the metric components are:
@@
Then, for such an energy density, the equation of motion of $b$ (67) becomes $\mathrm{d}_{r}\log b=-c_{1}$, which constrains the behavior of $c_{1}(r)$ for $r\to 0$ to $c_{1}(r)\sim r^{m}$ with $m\geq 1$. Similarly, the equation of motion for $a$ (66) implies that $c_{0}=\dot{a}-c_{1}a\sim r^{n}$ for $r\to 0$ with $n\geq 2$. Therefore, we obtain the effective Hamiltonian for regular metrics $H_{\mathrm{eff}}[a,b,r]=b(c_{0}(r)+c_{1}(r)a)$ with:
$\left|\begin{array}[]{cccc}c_{0}(r)&\underset{r\to 0}{\propto}&r^{n}&\textrm{with}\,\,n\geq 2\,,\vspace*{1mm}\\ c_{1}(r)&\underset{r\to 0}{\propto}&r^{m}&\textrm{with}\,\,m\geq 1\,,\end{array}\right.$ (107)
@@ VI.2 Improved KY metric
To this purpose, let us consider the simplest regular linear effective energy density, satisfying the conditions discussed above. As $c_{0}(r)$ is dimensionless and $c_{1}(r)$ has the dimension of the inverse of length, the smallest power laws satisfying the the conditions (107) reads:
@@
where we introduce Planck length factors to appropriately balance out the $r$-factors. The constant parameters $k_{0},k_{1}$ are dimensionless. Comparing to the effective energy density (97) for the KY metric which we derived in the previous section, we noticedw that this regular ansatz has the same structure except for a missing factor in $1/r^{2}$ and a constant term. We will further comment on this key difference below.
@@
[…]
@@
$\mathrm{d}s^{2}\underset{r\to 0}{\sim}$ $-\left(1-\frac{r^{2}}{6\sigma}\right)\frac{1}{4}e^{-\frac{r_{s}^{2}-r^{2}}{2\sigma}}dt^{2}$
@@
which is exactly the KY metric for $\eta=1$.
That is, the metric (VI.2) describes the whole interior region of the black hole $0\leq r\leq r_{s}$. It fits with the KY metric with $\eta=1$ in the semi-classical region $r\gg l_{p}$ and defines a fully regular completion of that metric to the central region $r\to 0$. It can therefore be considered as an improved version of the KY metric. Its dynamics contains the effects of quantum matter fields including quantum fluctuations consistent with the Weyl anomaly and excitations consistent with the entropy-area law, and possesses a mechanism that removes the singularity. It is remarkable that such a metric is derived naturally from the simplest effective Hamiltonian satisfying the regularity condition. This underlines the potential universality of this metric.
@@
$-{\cal T}{}_{t}{}^{t}={\cal T}{}_{r}{}^{r}=\frac{{\cal F}\left(\frac{r}{\sqrt{4\sigma}}\right)}{8\pi G\sqrt{\sigma}~{}r}=\left\{\begin{array}[]{l}\frac{1}{16\pi G\sigma}+\mathcal{O}(r^{2})\,{\rm{for}}\,r\to 0\,,\vspace*{2mm}\\ \frac{1}{8\pi Gr^{2}}+\mathcal{O}(r^{-4})\,{\rm{for}}\,r\gg l_{p}\,,\end{array}\right.$
@@
Figure 3: $-{\cal T}{}_{t}{}^{t}(r)={\cal T}{}_{r}{}^{r}(r)$ in (VI.2)(blue), ${\cal T}{}_{\theta}{}^{\theta}(r)$ in (122)(green), and (an extrapolation of) $\langle\Psi|-T{}_{t}{}^{t}(r)|\Psi\rangle_{KY}=\langle\Psi|T{}_{r}{}^{r}(r)|\Psi\rangle_{KY}|_{\eta=1}$ in (93) (orange). $18\pi G=$ 1, $\sigma=1$.
@@
$\textrm{with}\,\,\left|\begin{array}[]{l}C_{\alpha\beta\mu\nu}C^{\alpha\beta\mu\nu}=R_{\alpha\beta\mu\nu}R^{\alpha\beta\mu\nu}-2R_{\alpha\beta}R^{\alpha\beta}+\frac{1}{3}R^{2}\,,\vspace*{2mm}\\ \mathcal{G}=R_{\alpha\beta\mu\nu}R^{\alpha\beta\mu\nu}-4R_{\alpha\beta}R^{\alpha\beta}+R^{2}\,,\end{array}\right.$
@@
where we look for a suitable coupling constant $k$. In order to solve for $a(r)$ and $b(r)$ perturbatively, we expand the solutions to the equations of motion as:
@@ Appendix A Derivation of the Regularity conditions
We derive here the regularity conditions (105) and (106). We first consider the case for which
$\dot{a}\sim kr^{n},~{}~{}\frac{d}{dr}\log b\sim k^{\prime}r^{m}$ (132)
for $r\to 0$. Then, the Einstein tensors (IV.1) become
@@
Using the Einstein equations, the regularity for the energy density $-{\cal T}{}_{t}{}^{t}$ and radial pressure ${\cal T}{}_{r}{}^{r}$ thus requires

### v2

@@ V.3 Generating Kawai-Yokokura metrics
at leading orders for large mass, $a_{0}\gg l_{p}$. Note that we write here $r_{s}$ for the surface radius, instead of the Schwarzschild radius. Indeed, it must be underlined that $r_{s}>a_{0}$ and that there is no actual horizon [^8] Note that the proper length between $r=r_{s}$ and $r=a_{0}$ is of order $\sqrt{n}l_{p}\gg l_{p}$. . Phenomenologically, however, the exponentially large redshift of the KY metric (87) makes the imaging almost black and, seen from the exterior, the object behaves as a black hole in astrophysical cases with non-Planckian masses. The small deviation from classical Schwarzschild black holes is consistent with current observations CY .
@@
Indeed, it was shown in KY4 that the KY metric (87) is a non-perturbative solution in $\hbar$ for $1\leq\eta<2$ and $\sigma=\mathcal{O}(nl_{p}^{2})$, where $n$ is the number of the degrees of freedom that contribute to the entropy of the system. More precisely, for $n$ massless scalar fields minimally-coupled to gravity, we have KY4 :
@@
The KY metric leads to a couple of interesting physical insights. First, it can be obtained from several paths. The original way KMY is to study the time evolution of a spherical collapsing matter including the backreaction from the particle creation in the time-dependent spacetime according to (89). The final object at equilibrium in a heat bath is described by the KY metric [^9] There exist similar constructions incorporating various aspects of QFT fluctuations in the structure of black holes. A successful example are black stars that include the backreaction of vacuum polarization at every point inside the compact object (through a renormalized stress-energy tensor), creating an onion structure with Hawking radiation between each shell generating a negative pressure balancing out the gravitational pull Barcelo:2007yk ; Visser:2008rtf ; Barcelo:2009tpa . The dynamical structure of KY black holes is very similar, and further incorporates non-perturbative 4D quantum fluctuations (beyond the vacuum state) through a detailed balance of the 4D anomaly. The Hawking radiation within the compact object is a mere semi-classical expression of those quantum fluctuations. KY1 ; KY2 ; KY3 ; KY4 ; KY5 . Another path is to identify the KY metric as the entropy-maximizing configuration for a given surface area $\mathcal{A}$, saturating the Bousso bound Y2 .
@@
only for $l_{p}\ll r\leq r_{s}$, where the constant factor $b_{0}=\frac{\eta}{2}e^{-\frac{r_{s}^{2}}{4\eta\sigma}}$ encodes the mass (and thus the surface radius) of the black hole. Such an effective Hamiltonian is supposed to reflect the full non-perturbative coupling of quantum matter fields on the geometry (such as the curvature generated by the one-loop renormalized QFT corrections).
@@
We can check that solving the corresponding equations of motion for $a$ and $b$, (66) and (67), leads back to the desired behavior only for $r\gg l_{p}$:
@@
Solving exactly the equation of motion for $a$ leads to the desired behavior only for $r\gg l_{p}$,
@@
where $\mathrm{erfc}(x)\equiv 1-\frac{2}{\sqrt{\pi}}\int^{x}_{0}dye^{-y^{2}}$. Thus, we once again reproduce the KY metric, as expected. This illustrates the degeneracy in identifying effective Hamiltonians consistent with a given metric. Choosing a specific Hamiltonian over the other possibilities will have to be motivated by other considerations, such as a matching with modified gravity action principles or a direct derivation from a fundamental Lagrangian coupling renormalized matter fields to general relativity. [^10] Again, we can select a special one $c_{2}(r)$ by asking the consistency with (64) in the linear level for $r\gg l_{p}$.
@@
$\tilde{a}(r)=\left(C-\frac{2}{3}\int^{r}_{r_{0}}dr^{\prime}c_{3}(r^{\prime})e^{2\int^{r^{\prime}}_{r_{0}}dr^{\prime\prime}c_{1}(r^{\prime\prime})}\right)e^{-2\int^{r}_{r_{0}}dr^{\prime}c_{1}(r^{\prime})},$
where $C$ and $r_{0}$ are two integration constants [^11] The extra integration constant $C$ does not affect the behavior of the metric at $r\gg l_{P}$, but it suggests various versions and completions of the KY metric to the central region $0\leq r\lesssim l_{p}$. . Plugging the explicit expressions for $c_{1}$ and $c_{3}$ yields:
$\tilde{a}(r)$ $=Ce^{-\frac{r^{2}-r_{0}^{2}}{4\sigma\eta}}+\frac{1}{2\sigma\eta}\int^{r}_{r_{0}}dr^{\prime}\frac{1}{r^{\prime}}\left(1+\frac{4\sigma}{r^{\prime 2}}\right)e^{-\frac{r^{2}-r^{\prime 2}}{4\sigma\eta}}$
$\approx Ce^{-\frac{r^{2}-r_{0}^{2}}{4\sigma\eta}}+\frac{1}{2\sigma\eta}\frac{1}{r}\left(1+\frac{4\sigma}{r^{2}}\right)\int^{r}_{r_{0}}dr^{\prime}e^{-\frac{r(r-r^{\prime})}{2\sigma\eta}}$
@@
This shows that our type-$(+1)$ ansatz for effective Hamiltonians, linear in the lapse $b$ and polynomial in the mass function $a$, is rich enough to contain a whole family of Hamiltonians generating the KY metric for $r\gg l_{p}$. Note that the KY metric does not solve these Hamiltonians exactly but approximately for $r\gg l_{p}$, thus allowing flexibility for its completion in the Planckian region $0<r\lesssim l_{p}$. More precisely, the expression of the effective energy density $\rho_{\mathrm{eff}}(a,r)$ is not unique, and there are various expressions of $\rho_{\mathrm{eff}}(a,r)$ that reproduce the same KY metric for $r\gg l_{p}$ and describe different Planck-scale physics for $0\leq r\lesssim l_{p}$.
Then, one way to choose “natural” Hamiltonians from various admissible candidates is to consider the completion of the KY space-time to the whole black hole interior. As shown in the next section, this will allow us to identify a “simple” effective Hamiltonian that realizes a fully regular completion of the KY metric without singularity for the whole interior region $0\leq r\leq r_{s}$.
@@ VI.1 Conditions on the Energy Density
We first examine the conditions on $a(r)$ and $b(r)$ for the spherically-symmetric static metric (3) to be regular at $r=0$. For simplicity, we here consider a “typical” case for which
$\dot{a}\sim kr^{n},~~\frac{d}{dr}\log b\sim k^{\prime}r^{m}$ (105)
for $r\to 0$, although it ignores possible logarithmic or exponential behaviors, which occur in some scenarios. Checking the behavior of the curvatures as $r\to 0$, as proven in details in appendix A, one can show that the necessary behavior for the metric components are:
@@
Then, for such an energy density, the equation of motion of $b$ (67) becomes $\mathrm{d}_{r}\log b=-c_{1}$, which constrains the behavior of $c_{1}(r)$ for $r\to 0$ to $c_{1}(r)\sim r^{m}$ with $m\geq 1$. Similarly, the equation of motion for $a$ (66) implies that $c_{0}=\dot{a}-c_{1}a\sim r^{n}$ for $r\to 0$ with $n\geq 2$. Therefore, we obtain the effective Hamiltonian for regular metrics $H_{\mathrm{eff}}[a,b,r]=b(c_{0}(r)+c_{1}(r)a)/2G$ with:
[…]
@@ VI.2 Improved KY metric
To this purpose, let us consider the simplest regular linear effective energy density, [^12] We here focus on thermodynamically typical configurations and need $a$-dependence of $H_{\mathrm{eff}}$ (from (60)). Note that regular black-hole metrics in the literature 1968qtr..conf…87B ; Dymnikova:1992ux ; Ayon-Beato:1998hmi ; Dymnikova:2004zc ; Hayward:2005gi satisfy $\rho_{\mathrm{eff}}+p_{\mathrm{eff}}=0$ and are not typical. satisfying the conditions discussed above. As $c_{0}(r)$ is dimensionless and $c_{1}(r)$ has the dimension of the inverse of length, the smallest power laws satisfying the the conditions (108) reads:
@@
where we introduce Planck length factors to appropriately balance out the $r$-factors. The constant parameters $k_{0},k_{1}$ are dimensionless such that $k_{0}k_{1}>0$ for the configuration to be typical (from (60)). Comparing to the effective energy density (97) for the KY metric which we derived in the previous section, we notice that this regular ansatz has the same structure except for a missing factor in $1/r^{2}$ and a constant term. We will further comment on this key difference below.
@@
where ${\cal F}(x)\equiv e^{-x^{2}}\int^{x}_{0}dye^{y^{2}}$ is Dawson’s integral, and $C$ is an integration constant. This lapse factor $b(r)$ directly agrees with the KY metric (95) upon matching the parameters $(k_{0},k_{1})$ with the KY parameters $(\sigma,\eta)$ such that $2k_{0}k_{1}l_{P}^{2}=4\sigma\eta$. This makes sense dimension-wise since $\eta$ is dimensionless and $\sigma$ has the dimension of an area.
@@
$\mathrm{d}s^{2}\underset{r\to 0}{\sim}$ $-\left(1+\frac{r^{2}}{3\sigma}\right)\frac{1}{4}e^{-\frac{r_{s}^{2}}{2\sigma}}dt^{2}$
@@
which is exactly the KY metric for $\eta=1$. [^13] The improved KY metric (VI.2) becomes approximately the original KY metric (118) for $r\gg l_{p}$. Therefore, following the argument of Ref.Y1 , the position of the surface satisfying the Israel junction condition can be identified as (88), as long as the total mass $\frac{a_{0}}{2G}$ is large. That is, for the Schwarzschild external metric and the original KY internal metric, requiring that all components of the surface energy-momentum tensor be as small as possible (i.e., that the interior and exterior be as continuous as possible) leads to the surface location being chosen as (88).
That is, the metric (VI.2) describes the whole interior region of the black hole $0\leq r\leq r_{s}$. It fits with the KY metric with $\eta=1$ in the semi-classical region $r\gg l_{p}$ and defines a fully regular completion of that metric to the central region $r\to 0$. It can therefore be considered as an improved version of the KY metric. Its dynamics contains the effects of quantum matter fields including quantum fluctuations consistent with the Weyl anomaly and excitations consistent with the entropy-area law, and possesses a mechanism that removes the singularity. It is remarkable that such a metric is derived naturally from the simplest effective Hamiltonian satisfying the regularity condition and the typical condition. This underlines the potential universality of this metric.
@@
[…]
@@
[…]
@@
Another scenario?
The energy scale of the interior of the object described by the improved KY metric is close to the Planck scale, and it should be natural to assume that matter fields are massless and approximately conformal. As shown above, this led us to the mismatch of the two values of $\sigma$, meaning that the semi-classical Einstein equation is not valid at the central part.
Nevertheless, let us here try to suggest another scenario where the semi-classical dynamics is kept. Suppose that $N$ new massive scalar fields $\phi_{i}$ with Planck scale masses $m_{i}=\mathcal{O}(m_{p})$ appear around $r=0$ by some new physics (e.g. String ). Then, the right hand side in the trace part of the semi-classical Einstein equation (127) would gain additional terms $8\pi G\frac{1}{2}g^{\mu}{}_{\mu}\sum_{i=1}^{N}m_{i}^{2}\langle\phi_{i}^{2}\rangle$:
$\frac{1}{\sigma}$ $=\frac{32\pi l_{p}^{2}a_{W}}{3\sigma^{2}}-16\pi GNm_{p}^{2}C$
$\Rightarrow NC$ $=\frac{3}{128\pi^{2}c_{W}G}\left(\frac{4a_{W}}{c_{W}}-1\right),$ (128)
where we assume $\langle\phi_{i}^{2}\rangle\approx C={\rm const.}$ for $r\sim 0$ and set $\sigma=\sigma_{IR}$ (126). This would be another scenario within the semi-classical dynamics as long as the self-consistency is checked.
At the end of the day, more work on new dynamics around the center would be required to settle this issue.
@@
where we look for a suitable coupling constant kk. [^15] Note here that all terms in the bracket are dimensionless and should be proportional to $\frac{1}{\sigma}\sim\frac{1}{l_{p}^{2}}$, as expected as a non-perturbative effect. Therefore, the coefficient function of the quadratic term should be independent of $r$. In order to solve for $a(r)$ and $b(r)$ perturbatively, we expand the solutions to the equations of motion as:
@@ Appendix A Derivation of the Regularity conditions
We derive here the regularity conditions (106) and (107). We first consider the case (105):
$\dot{a}\sim kr^{n},~~\frac{d}{dr}\log b\sim k^{\prime}r^{m}~~{\rm for}~~r\to 0.$
Then, the Einstein tensors (IV.1) become
@@
Using the Einstein equations, the regularity for the energy density ${\cal T}{}_{t}{}^{t}$ and radial pressure ${\cal T}{}_{r}{}^{r}$ thus requires